OpenStudy (anonymous):

How many grams of calcium phosphate are theoretically produced if we start with 10.3 moles of Ca(NO3)2 and 7.6 moles of Li3PO4? Reaction: 3Ca(NO3)2 + 2Li3PO4 → 6LiNO3 + Ca3(PO4)2

OpenStudy (jfraser):

use each reactant separately, and find out how much calcium phosphate you could make in either situation. The lesser amount is the actual answer, because once one of the reactants runs out, you can't make anymore product

OpenStudy (abmon98):

Determine the limiting reagent which would be lithium phosphate ratio of moles of lithium phosphate to calcium phosphate 2:1 7.6/2=3.8 moles of calcium phosphate Number of moles=mass(g)/ molar mass (g/mol) 3.8*310=1178 grams